# Computing Lower and Upper Hitting Probabilities for Imprecise Markov Chains

### v1

@@ Abstract
We study the computation of lower and upper probabilities of hitting a target set of states for imprecise Markov chains. For these, transition uncertainty is modelled by a convex set of transition matrices. In the precise case, hitting probabilities are the minimal nonnegative solution of a linear system and admit a closed-form expression. We study the notion of reachability in the imprecise setting. The literature review highlights few different definitions of lower reachability; thus we explore the relations among them, presenting examples to clarify their logical implications. Using this revised definition of reachability for imprecise Markov chain, we partition the state space into classes of states whose hitting probabilities are trivially zero or one and those which require further computation. For these nontrivial states, we show that lower and upper hitting probabilities are the unique solutions of two nonlinear fixed-point equations. For the practical computation of lower and upper hitting probabilities, we propose iterative algorithms that alternate between solving a linear system and choosing an extreme point from the set of transition matrices. Numerical experiments demonstrate that, in practice, these algorithms converge in substantially fewer iterations than the theoretically established worst-case bound.
@@ 1 Introduction
Hitting probabilities play a central role in the analysis of stochastic systems, capturing the likelihood that a process enters a designated set of target states in some finite time. In the classical precise setting of homogeneous Markov chains, these probabilities arise as the minimal nonnegative solution of a linear system and admit a closed-form representation via the fundamental matrix [norrisbook1997markov, kemeny1976finite]. Hitting probabilities can also be regarded as discrete harmonic functions and enjoy a deep analogy with node voltages in potential theory and electrical networks [DoyleSnell1984].
@@
In practical applications there is often uncertainty about the transition dynamics, which can be estimated from limited data or elicited from domain experts, yielding a convex set of transition matrices rather than a single, exact one. A natural and rigorous way to represent this kind of epistemic uncertainty is through credal sets – convex sets of probability distributions – leading to the framework of imprecise Markov Chains (IMCs) [DeCooman2009, HermansSkulj2014, hartfiel_seneta_1994]. In this framework, each row of the transition matrix is specified by a convex set of admissible probability distributions, and the inference quantities of interest are lower and upper hitting probabilities. These bound the true probability of reaching the target states under all compatible transition matrices. These values provide conservative guarantees that are crucial in high-stakes decision-making contexts. They naturally connect to robust control, since they provide worst-case bounds used in min–max robust MDPs under transition uncertainty [NilimElGhaoui2005], and to stochastic-game formulations, as they correspond to value bounds in zero-sum reachability games [Bovy2024].
@@
Our contributions build on the work of Krak et al. [krak2019hitting, krak2021comphit]. They developed characterisations of hitting times and probabilities for IMCs under various model semantics and presented algorithms for computing lower and upper hitting times. We extend this work by (i) discussing and clarifying the definition of reachability in the imprecise setting, (ii) giving a more general characterisation of lower and upper hitting probabilities and (iii) proposing iterative algorithms to compute these quantities.
The structure of the paper is as follows. Section 2 introduces the preliminary notions needed for the following sections. In particular, we define hitting probabilities for homogeneous Markov chains, we explain the concept of imprecise Markov chains, and we introduce lower and upper transition operators. Section 3 begins our contributions. There, we formalise the notions of lower and upper reachability in the imprecise setting; we review the different definitions that appear in the literature, clarify the logical relationships between them, and give a few illustrative examples. With these definitions in hand, we isolate those states whose (lower/upper) hitting probabilities are trivial, either zero or one. We then prove that for nontrivial states, lower and upper hitting probabilities are the unique solutions of two nonlinear fixed-point equations.
@@
In Section 4, we turn our attention to developing iterative algorithms to compute lower and upper hitting probabilities. Inspired by the work of Krak et al. [krak2021comphit] on hitting times, we propose iterative algorithms to compute lower and upper hitting probabilities that alternate between solving a linear system (restricted to a suitable subspace) and selecting an extreme point transition matrix from the credal sets. We rigorously prove their convergence and, in Section 5, we provide a detailed analysis of their computational cost. We conclude the paper with Section 6 which is dedicated to numerical experiments. Here we show how the theoretical bound on the number of iterations previously established is never attained in practice, and we point out an interesting relation between the density of the graph and the average number of iterations needed for convergence.
@@ 2 Preliminaries
This section introduces the basic concepts of stochastic processes [Dudley2002], Markov chains [norrisbook1997markov, kemeny1976finite], hitting probabilities for homogeneous Markov chains [norrisbook1997markov], and imprecise Markov chains [DeCooman2009, HermansSkulj2014, hartfiel_seneta_1994].
@@ 2.1 Stochastic Processes and Markov Chains
where we set $X_{n}(\omega)=\omega(n)$ for all $\omega\in\Omega$ and all $n\in\mathbb{N}_{0}$ [Dudley2002]. Let $\mathbb{P}$ be a probability measure on the measurable space $(\Omega,\mathcal{F})$, so that $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. The probability measure $\mathbb{P}$ determines the law of the stochastic process; in the sequel, we will identify the process with the probability measure $\mathbb{P}$, which allows us to consider different processes $\mathbb{P}$ over the same measurable space $(\Omega,\mathcal{F})$ [krak2019hitting].
A stochastic process $\mathbb{P}$ is said to be a Markov chain if it satisfies the Markov property:
@@
for all $x_{0},\dots,x_{n},x_{n+1}\in\mathcal{X}$ and all $n\in\mathbb{N}_{0}$. This property is often interpreted as the process being memoryless; the transition probabilities depend only on the current state, and any previous history is not relevant. A Markov chain $\mathbb{P}$ is said to be (time-)homogeneous if
$\mathbb{P}(X_{n+1}=x\mid X_{n}=y)=\mathbb{P}(X_{1}=x\mid X_{0}=y)$ (2)
@@
$\mathbb{P}_{T}(X_{n+1}=x\mid X_{n}=y)=T{(x,y)}$ (3)
for all $x,y\in\mathcal{X}$ and for all $n\in\mathbb{N}_{0}$. Thus, transition matrices are often used as the canonical parameter to specify homogeneous Markov chains.
Definition 2.1 (Reachability [norrisbook1997markov]).
Let $\in\mathcal{X}, xyx\neq y$, and let $T$ be a transition matrix. We say that $x$ reaches $y$, and we denote it with $\to^{T}$”, if there exists $n\in\mathbb{N}$ and a sequence of states $x=x_{0},\ldots,x_{n}=y$ in $\mathcal{X}$ such that $T(x_{k},x_{k+1})>0$ for all $0\leq k\leq n-1$, and $x_{k}\neq x_{j}$ for all $0\leq k<j\leq n$.
@@
Proposition 2.2 ([norrisbook1997markov, Theorem 1.2.1]).
@@
    $x\to^{T}y$,
@@ 2.2 Hitting Probabilities for Markov Chains
This quantity is a Boolean function having value one if the path reaches $A$ and zero otherwise.
@@
The hitting probability of a process $\mathbb{P}$ starting in $x\in\mathcal{X}$ can also be seen as
$\mathbb{P}\,\Big(\exists n\in\mathbb{N}:X_{n}\in A\mid X_{0}=x\Big)=\mathbb{P}\,\Bigg(\bigcup_{n\in\mathbb{N}_{0}}\{X_{n}\in A\}\ \Big|\ X_{0}=x\Bigg),$
i.e. the probability that $A$ is reached in some number of steps. In the remainder of the paper, we collect these hitting probabilities in the real-valued function $p_{A}^{\mathbb{P}}:\mathcal{X}\to[0,1]$ defined by
@@
$=$ $\{x\in A^{c}\mid\forall y\in A,\forall n\in\mathbb{N}_{0},:T^{n}(x,y)=0\}.$
We can characterise the hitting probabilities of a homogeneous Markov chain as the minimal nonnegative solution to a linear system of equations.
Proposition 2.5 ([norrisbook1997markov, Theorem 1.3.2]).
Let $\mathbb{P}_{T}$ be a homogeneous Markov chain defined on the state space $\mathcal{X}$ with transition matrix $T$. Then $p^{T}(x)=1$ for all $x\in A$, and $p^{T}(x)=0$ if and only if $x\in\mathcal{C}_{T}$. Moreover, $p^{T}$ is the (unique) minimal nonnegative solution of the system
@@ 2.3 Restrictions and Extensions of Functions and Matrices to a Subspace
This extension is the right inverse of restriction, namely
@@
for all $g\in\mathbb{R}^{\mathcal{S}}$.
Let $M:\mathbb{R}^{\mathcal{Y}}\to\mathbb{R}^{\mathcal{Y}}$ . Define the restriction of $M$ to $\mathcal{S}$ as
@@ 2.4 Closed-Form Solution for Hitting Probabilities
In this section, we derive a closed-form expression for hitting probabilities for homogeneous Markov chains. Although most of the material in this section appears in standard references (e.g. [norrisbook1997markov, kemeny1976finite, DoyleSnell1984, GrinsteadSnelll1997]), we present it here in full detail, since the terminology and results will be essential for the subsequent development in the imprecise setting.
By Proposition 2.5, a homogeneous Markov chain $\mathbb{P}_{T}$ with transition matrix $T$, hitting probabilities are trivial for states in $A\cup\mathcal{C}_{T}$. The system of equations in (4), as we restrict to $A^{c}\setminus\mathcal{C}_{T}$, can be reduced to
@@
Hence, if we can show that the matrix $(I-T|_{A^{c}\setminus\mathcal{C}_{T}})$ is invertible, we can uniquely characterise $p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}$ and have a closed-form expression to compute it. The structure of the argument is similar to the reasoning used in Krak [krak2021comphit, Lemma 10].
We proceed with the following lemma, the proof of which can be found in the appendix.
@@
Let $T$ be a transition matrix. Fix any $x\in A^{c}\setminus\mathcal{C}_{T}$ and $n\in\mathbb{N}$ with $n>1$. Then if $T^{n}\mathbb{1}_{A}](x)>0$ and $T^{m}\mathbb{1}_{A}](x)=0$ for all $m\in\mathbb{N}$ with $m<n$, there exists some $y\in A^{c}\setminus\mathcal{C}_{T}$, $y\neq x$ such that $T(x,y)>0$ and $T^{n-1}\mathbb{1}_{A}](y)>0$.
Let $\mathbf{0}$ and $\mathbf{1}$ be the vectors made of all zeros and ones [^3] The size of these vectors will be clear for the context., respectively, and let $\mathcal{S}\subseteq\mathcal{X}$. Then, the quantity $(T|_{\mathcal{S}})^{k}\mathbf{1}](x)$, for $x\in\mathcal{S}$ and $k\in\mathbb{N}_{0}$, measures the probability that a homogeneous Markov chain with transition matrix $T$ starting in state $x$ stays in $\mathcal{S}$ after $k$ steps. Hence, it is not very surprising that this quantity is bounded below by zero and above by one.
@@
Let $T$ be a transition matrix and consider an arbitrary subset of states $\mathcal{S}\subseteq\mathcal{X}$. Let $f,g\in\mathbb{R}^{\mathcal{S}}$ be such $f\leq g$. Then $(T|_{\mathcal{S}})^{k}f\leq(T|_{\mathcal{S}})^{k}g$ for all $k\in\mathbb{N}_{0}$. Moreover, we have that
@@
Let $T$ be a transition matrix, and let $A$ be the target set. Then there exists $n\in\mathbb{N}$ such that for all $k\in\mathbb{N}$ with $k\geq n$, it holds that
@@
Since the intuition for Lemma 2.9 was that it holds because every state in $A^{c}\setminus\mathcal{C}_{T}$ will reach a state outside of $A^{c}\setminus\mathcal{C}_{T}$, the result actually holds for every subset of $A^{c}\setminus\mathcal{C}_{T}$ as well. This slightly more general result will prove useful in the imprecise setting later on.
@@
Let $T$ be a transition matrix. Consider an arbitrary subset of states $\mathcal{S}\subset\mathcal{X}$ such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then there exists $n\in\mathbb{N}$ such that for all $k\in\mathbb{N}$ with $k\geq n$, it holds that
@@
Using this result, we can conclude that powers of the transition matrix $T$ restricted to $A^{c}\setminus\mathcal{C}_{T}$ are smaller in norm than 1.
@@
Let $T$ be a transition matrix. Consider an arbitrary subset of states $\mathcal{S}\subset\mathcal{X}$ such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then there exists some $n\in\mathbb{N}$ such that
@@
Let $T$ be a transition matrix. Consider an arbitrary subset of states $\mathcal{S}\subset\mathcal{X}$ such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then
@@
By Corollary 2.11, there exists some $n\in\mathbb{N}$ such that $||(T|_{A^{c}\setminus\mathcal{S}})^{n}||<1$. This means that the restriction of $T$ to $A^{c}\setminus\mathcal{S}$ eventually becomes a contraction and the probability of remaining there decays to zero. Hence, the Neumann series $\sum_{k=0}^{\infty}(T|_{A^{c}\setminus\mathcal{S}})^{k}$ converges and the result follows (Kemeny and Snell [kemeny1976finite] give more details). ∎
The next proposition gives a closed-form expression for the vector of hitting probabilities of a given transition matrix $T$.
@@
Let $\mathbb{P}_{T}$ be a homogeneous Markov chain defined on the state space $\mathcal{X}$ with transition matrix $T$. Then $p^{T}(x)=1$ for $x\in A$ and $p^{T}(x)=0$ if and only if $x\in\mathcal{C}_{T}$. Furthermore, the vector $p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}$ is given by
$p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=\bigl(I-T|_{A^{c}\setminus\mathcal{C}_{T}}\bigr)^{-1}(T\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{C}_{T}}.$ (8)
@@
where we applied Lemma 2.12 to guarantee the existence of the inverse. ∎
The last lemma we present in this section is required for results in the imprecise setting later on. Given a transition matrix $T$, consider the induced directed weighted graph containing as vertices the states of $\mathcal{X}$ and directed edges with weight $T(x,y)$ for all $x,y\in\mathcal{X}$. Then, the following lemma establishes that, from any state $x\in\mathcal{X}$ which can reach $A$, there must be a (directed) simple path from $x$ to $A$ using only nonzero weight edges for which the hitting probabilities of the states along the way are nondecreasing as we approach $A$. Essentially, this says that from every such state $x$, there must be a way to “get closer” (that is, increase the hitting probability) to $A$.

### v2

@@ Abstract
We study the computation of lower and upper probabilities of hitting a target set of states for imprecise Markov chains, where transition uncertainty is modelled by a convex set of transition matrices. In the precise case, hitting probabilities are the minimal nonnegative solution of a linear system and admit a closed-form expression. We investigate the notion of reachability in the imprecise setting. The literature review highlights several different definitions of lower reachability; thus, we explore the relations among them and present examples to clarify their logical implications. Using this revised definition of reachability for imprecise Markov chains, we partition the state space into classes of states whose hitting probabilities are trivially zero or one, and those which require further computation. For these nontrivial states, we show that the lower hitting probability is the unique solution of a nonlinear fixed-point equation, while the same does not hold for upper hitting probabilities. For the practical computation of lower and upper hitting probabilities, we propose iterative algorithms that alternate between solving a linear system and choosing an extreme point from the set of transition matrices. Numerical experiments demonstrate that, in practice, these algorithms converge in substantially fewer iterations than the theoretically established worst-case bound.
@@ 1 Introduction
Hitting probabilities play a central role in the analysis of stochastic systems, capturing the likelihood that a process enters a designated set of target states in some finite time. In the classical precise setting of homogeneous Markov chains, these probabilities arise as the minimal nonnegative solution of a linear system and admit a closed-form representation via the fundamental matrix [21, 18]. Hitting probabilities can also be regarded as discrete harmonic functions and enjoy a deep analogy with node voltages in potential theory and electrical networks [8].
@@
In practical applications there is often uncertainty about the transition dynamics, which can be estimated from limited data or elicited from domain experts, yielding a convex set of transition matrices rather than a single, exact one. A natural and rigorous way to represent this kind of epistemic uncertainty is through credal sets – convex sets of probability distributions – leading to the framework of imprecise Markov chains (IMCs) [5, 14, 12]. In this framework, each row of the transition matrix is specified by a convex set of admissible probability distributions, and the inference quantities of interest are lower and upper hitting probabilities. These bound the true probability of reaching the target states under all compatible transition matrices. These values provide conservative guarantees that are crucial in high-stakes decision-making contexts. They naturally connect to MDPs as they correspond to the optimal value in reachability problems [23, 2], and to stochastic-game formulations as they correspond to value bounds in zero-sum reachability games [3].
@@
Our contributions build on the work of Krak et al. [19, 20]. They developed characterisations of hitting times and probabilities for IMCs under various model semantics and presented algorithms for computing lower and upper hitting times. We extend this work by (i) discussing and clarifying the definition of reachability in the imprecise setting, (ii) giving a more general characterisation of lower and upper hitting probabilities, and (iii) proposing iterative algorithms to compute these quantities.
@@
The structure of the paper is as follows. Section 2 introduces the preliminary notions needed for the following sections. In particular, we define hitting probabilities for homogeneous Markov chains, we explain the concept of imprecise Markov chains, and we introduce lower and upper transition operators. Section 3 begins our contributions. There, we formalise the notions of lower and upper reachability in the imprecise setting; we review the different definitions that appear in the literature, clarify the logical relationships between them, and give a few illustrative examples. With these definitions in hand, we isolate those states whose (lower/upper) hitting probabilities are trivial, either zero or one. We then prove that for nontrivial states, the lower hitting probability is the unique solution of a nonlinear fixed-point equation. Krak et al. [19] characterise the upper hitting probability as the minimal nonnegative solution of a nonlinear fixed-point equation, but this solution need not be unique. We illustrate this phenomenon with a counterexample and then show that uniqueness is restored under stronger assumptions on the set $\mathcal{T}$. In Section 4, we turn our attention to developing iterative algorithms to compute lower and upper hitting probabilities. Inspired by the work of Krak [20] on hitting times, we propose iterative algorithms to compute lower and upper hitting probabilities that alternate between solving a linear system (restricted to a suitable subspace) and selecting an extreme point transition matrix from the credal sets. We rigorously prove their convergence and, in Section 5, we provide a detailed analysis of their computational cost. The empirical performance of these algorithms is explored in Section 6, which is dedicated to numerical experiments. Here we show how the theoretical bound on the number of iterations previously established is never attained in practice, and we point out an interesting relation between the density of the graph and the average number of iterations needed for convergence. We conclude with some summarising remarks and pointers to possible future work in Section 7.
@@ 2 Preliminaries
This section introduces the basic concepts of stochastic processes [9], Markov chains [21, 18], hitting probabilities for homogeneous Markov chains [21], and imprecise Markov chains [5, 14, 12].
@@ 2.1 Stochastic Processes and Markov Chains
where we set $X_{n}(\omega)=\omega(n)$ for all $\omega\in\Omega$ and all $n\in\mathbb{N}_{0}$ [9]. Let $\mathbb{P}$ be a probability measure on the measurable space $(\Omega,\mathcal{F})$, so that $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. The probability measure $\mathbb{P}$ determines the law of the stochastic process; in the sequel, we identify the process with the probability measure $\mathbb{P}$, which allows us to consider different processes $\mathbb{P}$ over the same measurable space $(\Omega,\mathcal{F})$ [19].
A stochastic process $(X_{n})$ is said to be a Markov chain if it satisfies the Markov property:
@@
for all $x_{0},\dots,x_{n},x_{n+1}\in\mathcal{X}$ and all $n\in\mathbb{N}_{0}$. This property is often interpreted as the process being memoryless; the transition probabilities depend only on the current state, and any previous history is not relevant. A Markov chain $(X_{n})$ is said to be (time-)homogeneous if
$\mathbb{P}(X_{n+1}=y\mid X_{n}=x)=\mathbb{P}(X_{1}=y\mid X_{0}=x)$ (2)
@@
$\mathbb{P}_{T}(X_{n+1}=y\mid X_{n}=x)=T{(x,y)}$ (3)
for all $x,y\in\mathcal{X}$ and all $n\in\mathbb{N}_{0}$. Thus, transition matrices are often used as the canonical parameter to specify homogeneous Markov chains.
Definition 2.1 ([21, Section 1.2]).
Let $\in\mathcal{X}, xyx\neq y$, and let $T$ be a transition matrix. We say that $x$ reaches $y$, denoted by $x\to^{T}y$”, if
$\mathbb{P}_{T}(X_{n}=y\text{ for some $n\in\mathbb{N}_{0}$}\mid X_{0}=x)>0.$
@@
Proposition 2.2 ([21, Theorem 1.2.1]).
@@
    there exists $n\in\mathbb{N}$ and a sequence of states $x=x_{0},\ldots,x_{n}=y$ in $\mathcal{X}$ such that $T(x_{k},x_{k+1})>0$ for all $0\leq k\leq n-1$, and $x_{k}\neq x_{j}$ for all $0\leq k<j\leq n$,
@@ 2.2 Hitting Probabilities for Markov Chains
This quantity is a Boolean function that has a value of one if the path reaches $A$ and zero otherwise.
@@
The hitting probability of a process $\mathbb{P}$ starting in $x\in\mathcal{X}$ can also be written as $\mathbb{P}(X_{n}\in A\text{ for some $n\in\mathbb{N}_{0}$}\mid X_{0}=x)$, i.e. the probability that $A$ is reached in some number of steps. In the remainder of the paper, we collect these hitting probabilities in the real-valued function $p_{A}^{\mathbb{P}}:\mathcal{X}\to[0,1]$ defined as
@@
$=$ $\{x\in A^{c}\mid\forall y\in A,\forall n\in\mathbb{N}_{0}:T^{n}(x,y)=0\}.$
We can characterise the hitting probabilities of a homogeneous Markov chain as the minimal nonnegative solution of a linear system of equations.
Proposition 2.5 ([21, Theorem 1.3.2]).
Let $\mathbb{P}_{T}$ be a homogeneous Markov chain with transition matrix $T$. Then $p^{T}(x)=1$ for all $x\in A$ and $p^{T}(x)=0$ if and only if $x\in\mathcal{C}_{T}$. Moreover, $p^{T}$ is the minimal nonnegative solution of the system
@@
Proof.
The only claim not explicitly given in Norris [21, Theorem 1.3.2] is that $p^{T}(x)=0$ if and only if $x\in\mathcal{C}_{T}$. This follows directly from the definition of $\mathcal{C}_{T}$ and Proposition $2.$ ∎
@@ 2.3 Restrictions and Extensions of Functions and Matrices to a Subspace
This extension is the right inverse of the restriction, namely
@@
for all $g\in\mathbb{R}^{\mathcal{S}}$. Let $M:\mathbb{R}^{\mathcal{Y}}\to\mathbb{R}^{\mathcal{Y}}$ . Define the restriction of $M$ to $\mathcal{S}$ (or, rather, to $\mathbb{R}^{\mathcal{S}}$) as
@@ 2.4 Closed-Form Solution for Hitting Probabilities
In this section, we derive a closed-form expression for hitting probabilities for homogeneous Markov chains. Most of the material in this section appears in standard references [e.g 21, 18, 8, 11], however we present it here in full detail, since the terminology and results are essential for the subsequent development in the imprecise setting.
By Proposition 2.5, hitting probabilities of a homogeneous Markov chain $\mathbb{P}_{T}$ with transition matrix $T$ are trivial for states in $A\cup\mathcal{C}_{T}$. The system of equations in (4), as we restrict to $A^{c}\setminus\mathcal{C}_{T}$, can be reduced to
@@
Hence, if we can show that the matrix $(I-T|_{A^{c}\setminus\mathcal{C}_{T}})$ is invertible, we can characterise $p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}$ and have a closed-form expression to compute it.
We proceed with the following lemma, whose proof can be found in the appendix.
@@
Let $T$ be a transition matrix. Fix any $x\in A^{c}\setminus\mathcal{C}_{T}$ and $n\in\mathbb{N}$ with $n\geq 2$. If $T^{n}\mathbb{1}_{A}](x)>0$ and $T^{m}\mathbb{1}_{A}](x)=0$ for all $m<n$, then there exists some $y\in A^{c}\setminus\mathcal{C}_{T}$, $y\neq x$ such that $T(x,y)>0$ and $T^{n-1}\mathbb{1}_{A}](y)>0$.
Let $\mathbf{0}$ and $\mathbf{1}$ be the vectors made of all zeros and ones [^3] The size of these vectors will be clear from the context., respectively, and let $\mathcal{S}\subseteq\mathcal{X}$. Then, the quantity $(T|_{\mathcal{S}})^{k}\mathbf{1}](x)$, for $x\in\mathcal{S}$ and $k\in\mathbb{N}_{0}$, is equal to
$[(T|_{\mathcal{S}})^{k}\mathbf{1}](x)$ $=\sum_{y\in\mathcal{S}}(T|_{\mathcal{S}})^{k}(x,y)$
$=\sum_{y\in\mathcal{S}}\sum_{z_{1},\dots,z_{k-1}\in\mathcal{S}}T|_{\mathcal{S}}(x,z_{1})\cdots T|_{S}(z_{k-1},y)$
$=\mathbb{P}_{T}(X_{1}\in\mathcal{S},\dots,X_{k}\in\mathcal{S}\mid X_{0}=x),$
i.e. the probability that a homogeneous Markov chain with transition matrix $T$ starting in state $x$ stays in $\mathcal{S}$ after $k$ steps. Hence the following lemma holds.
@@
Let $T$ be a transition matrix and let $\mathcal{S}\subseteq\mathcal{X}$. Let $f,g\in\mathbb{R}^{\mathcal{S}}$ be such that $f\leq g$. Then $(T|_{\mathcal{S}})^{k}f\leq(T|_{\mathcal{S}})^{k}g$ for all $k\in\mathbb{N}_{0}$. Moreover, we have that
@@
Let $T$ be a transition matrix and let $A$ be the target set. There exists $n\in\mathbb{N}$ such that for all $k\in\mathbb{N}$ with $k\geq n$, it holds that
@@
Since the intuition for Lemma 2.9 was that it holds because every state in $A^{c}\setminus\mathcal{C}_{T}$ reaches a state outside of $A^{c}\setminus\mathcal{C}_{T}$, the result actually holds for every subset of $A^{c}\setminus\mathcal{C}_{T}$ as well. This slightly more general result will prove useful in the imprecise setting later on.
@@
Let $T$ be a transition matrix and let $\mathcal{S}\subset\mathcal{X}$ be such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then there exists $n\in\mathbb{N}$ such that for all $k\in\mathbb{N}$ with $k\geq n$, it holds that
@@
Using this result, we can conclude that powers of the transition matrix $T$ restricted to $A^{c}\setminus\mathcal{S}$ are smaller than one in norm.
@@
Let $T$ be a transition matrix and let $\mathcal{S}\subset\mathcal{X}$ be such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then there exists some $n\in\mathbb{N}$ such that
@@
Let $T$ be a transition matrix and let $\mathcal{S}\subset\mathcal{X}$ be such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then
@@
By Corollary 2.11, there exists some $n\in\mathbb{N}$ such that $||(T|_{A^{c}\setminus\mathcal{S}})^{n}||<1$. This means that the restriction of $T$ to $A^{c}\setminus\mathcal{S}$ eventually becomes a contraction and the probability of remaining there decays to zero. Hence, the Neumann series $\sum_{k=0}^{\infty}(T|_{A^{c}\setminus\mathcal{S}})^{k}$ converges and the result follows (Kemeny and Snell [18, Theorem $1.1$]). ∎
The next proposition gives a closed-form expression for the vector of hitting probabilities of $\mathbb{P}_{T}$.
@@
Let $\mathbb{P}_{T}$ be a homogeneous Markov chain with transition matrix $T$. Then $p^{T}(x)=1$ for all $x\in A$ and $p^{T}(x)=0$ if and only if $x\in\mathcal{C}_{T}$. The vector $p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}$ is given by
$p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=\bigl(I-T|_{A^{c}\setminus\mathcal{C}_{T}}\bigr)^{-1}(T\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{C}_{T}},$ (8)
and is the unique solution to
$p^{T}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathcal{C}_{T}}\cdot Tp^{T}.$ (9)
@@
where we applied Lemma 2.12 to guarantee the existence of the inverse. It then follows that $p^{T}$ is the unique solution to (9) ∎
The last lemma we present in this section is required for results in the imprecise setting later on. Given a transition matrix $T$, consider the induced directed weighted graph containing as vertices the states of $\mathcal{X}$ and directed edges with weight $T(x,y)$ for all $x,y\in\mathcal{X}$. Then, the following lemma establishes that, from any state $x\in\mathcal{X}$ which can reach $A$, there must be a (directed) simple path from $x$ to $A$ using only nonzero weight edges for which the hitting probabilities of the states along the way are nondecreasing as we approach $A$. Essentially, this says that from every state $x\in A^{c}\setminus\mathcal{C}_{T}$, there must be a way to “get closer” (i.e. increase the hitting probability) to $A$.